I need help on a proportional problem that I seen on the internet and I need it to be explained to me in lamen terms EXAMPLE:12/24=X/100 how do you determine the answer for the variable "x"?
Answer by xcentaur(357)

12/24=X/100
12 - X
24 - 100
Consider them as such:
12 is to X what 24 is to 100
Cross multiply to get the value of X
X=(12*100)/24=50


In other words,the ratio of 12 to 24 is the same ratio of X to 100,that is half.
If the ratio was to change,the value of X would accordingly change also.


Hope this helps,
best of luck.